OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 30%. What percent of the original volume is the volume of the new prism?

jimthompson5910 (jim_thompson5910):

hint: each dimension is multiplied by 1.3

OpenStudy (anonymous):

B*H*L=V 1.3B*1.3H*1.3L= 1.3^3V

OpenStudy (souvik):

volume increased by 1.2%

OpenStudy (whpalmer4):

@souvik how could that possibly be correct? you increase all the dimensions by 30%, and the volume only changes by 1.2%?

OpenStudy (whpalmer4):

Take a cube. Increase just one dimension (the height, for example) by 30%, what happens to the volume? It increases by 30%. How could increasing all three dimensions result in an increase of just over 1%?

OpenStudy (souvik):

new volume 1.3^3V volume increased= 1.3^3V-1^3V=1.197V~1.2V

OpenStudy (whpalmer4):

Agreed, but that's not an increase of 1.2%.

OpenStudy (anonymous):

What PERCENT of the original volume is the volume of the new prism? all you need to do is 1.3*1.3*1.3

OpenStudy (souvik):

oops in percentage it is 120%

OpenStudy (whpalmer4):

right. it's a 120% increase, but as @ConDawg points out, we aren't asked to find the increase, we're asked to find the relative size. It's 2.197 V or 219.7% of V

OpenStudy (souvik):

yeah right...@whpalmer4